when you have MC = 0 & P = abQ how to compute consumer surplus, producer surplus and deadweight loss due to their monopoly power?
Answer #1

P= a-bQ TR = P Q = aQ - bO² MR = a - 260 MR=mc a- 2bQ = 0 2bQ = a - Q = a 2b 야 ab을을 을 2b (44) که ر ماده ی جهان Produces suspl